If astronauts could travel at v=0.950 c, we on Earth would say it takes 4.42 years to reach Alpha Centauri. The astronauts disagree.

 How much time passes on the astronauts' clocks?

c = 3×108

1
Expert's answer
2022-04-22T16:43:16-0400

"t=t_0\sqrt{1-\frac{v^2}{c^2}}=1.38" years.